Obs # Age Obs # Age Obs # Age Obs # Age Obs # Age 1...

Obs #

Age

Obs #

Age

Obs #

Age

Obs #

Age

Obs #

Age

1

2019

11

2019

21

1976

31

2019

41

2006

2

1998

12

2019

22

2013

32

2018

42

2013

3

2019

13

2019

23

2019

33

2019

43

1982

4

1995

14

1980

24

1994

34

1997

44

2019

5

2018

15

2019

25

1979

35

2015

45

1988

6

2011

16

2016

26

1974

36

2019

46

2019

7

1974

17

1998

27

1984

37

2019

47

1999

8

1999

18

2018

28

2017

38

2019

48

2019

9

2017

19

2017

29

2019

39

2019

49

2001

10

2012

20

2015

30

2019

40

1986

50

2005

Use your sample to compute a 98% confidence interval for the mean age of pennies in circulation. Show all necessary calculations including the standard error and how you found the critical value.
Answer #1

mean=   2,007.14    [Excel function used -> AVERAGE]
sd=   14.934    [Excel function used -> STDEV]
n=   50.00    [Excel function used -> COUNT]
alpha=   2%  

critical value, z(a/2) = z(0.02/2) = (NORMSINV(1-0.02/2) = 2.326

SE = (sd/sqrt(n)) = (14.9339/sqrt(50)) = 2.111972

CI = mean +- z(a/2,n-1)*SE   
lower   = 2007.14 - 2.326*2.111972 = 2002.23
upper   = 2007.14 + 2.326*2.111972 = 2012.05

i am 98% confident that estimated mean age of pennies in circulation lie in the interval (2002.23, 2012.05).